The ages of a certain group of students in a junior high school follow a normal

distribution with a mean of 13 and a standard deviation of 1.5. Find the probability

that a randomly selected student is between 10-16 years old.


1
Expert's answer
2022-04-06T10:08:10-0400

We have a normal distribution, "\mu=13, \sigma=1.5."

Let's convert it to the standard normal distribution, "z=\cfrac{x-\mu}{\sigma};"

"z_1=\cfrac{10-13}{1.5}=-2, \\z_2=\cfrac{16-13}{1.5}=2,"


"P(10<X<16)=P(-2<Z<2)=\\=P(Z<1.5)-P(Z<-1.5)="

"=0.9332-0.0668=0.8664" ​(from z-table).